A gas with an initial volume of 24.0 L at a pressure of 565 mmHg is compressed by a pressure increase of 0.25 Pa. What is the final volume (L) of the gas, assuming the temperature does not change?
1
Expert's answer
2013-03-13T09:02:13-0400
This question can be solved using the ideal gas law:

PV = nRT

But even if you have two different conditions the amount is the same (R and T is constants), so:

0.1333 kPa in 1 mmHg
75.3145 kPa in 565 mmHg

P1V1 = nRT1
n = P1V1/RT1
n = P2V2/RT2
P1V1/RT1 = P2V2/RT2
P1V1 = P2V2
V2 = P1V1/P2

V2 = 75.3145 * 24.0 L / 75.31475 = 24.0